The reason for the warming-up of NOTCam that in the end caused the problems with the detector at the beginning of the semester (see above and the previous report) was that the motor controller of NOTCam had not been switched on after dismounting from the telescope thereby giving null input to the temperature logs, something that was not ``detected'' by our temperature monitoring system as a fault. The system has now been improved and it now checks if there are new readings twice per day and if not an alarm will be sent by email to the staff.
We still have some minor problems related to filling of LN - slight vacuum leaks which are however rapidly recovered by the cryostat - which are understood to be related to thermal stress around the fill-tube o-rings. A new type of LN filling nozzle was made drilling a cylinder from a specially ordered highly insulating material. The new nozzle is longer than the old one and occupies the whole width of the fill-tube, thereby avoiding leaks to the outer part where the sensitive o-rings are, and at the same time this material insulates between the LN flow and the inner metal of the fill-tube. The experience using the new nozzle is very good, and it is now easier to avoid spilling than before. There are still some very minor changes in the vacuum upon every filling, but quite small (about 610 mbar) and the cryostat recovers quickly. The vacuum 5 months after closing/pumping is better now (110 mbar) than one year ago (510 mbar).
Thomas Augusteijn 2010-02-09